# VeFA: Vector-Based Feature Space Adaptation for Mitigating Forgetting in Fine-Tuning

### v3

@@ Abstract
Catastrophic forgetting is a well-documented challenge in model fine-tuning, particularly under small-data downstream regimes. Existing parameter-efficient fine-tuning methods operate primarily in weight space, which can cause over-specialization of the model to the limited observations of the downstream data. Recent work reveals that one mechanism underlying such forgetting is the introduction of intruder dimensions into the representation space during fine-tuning. To mitigate overwriting of pre-trained knowledge and enhance robustness, we propose Vector-based Feature Space Adaptation (VeFA), a new fine-tuning scheme that operates directly in feature space and avoids generating intruder dimensions. VeFA applies element-wise transformation to individual feature channels, ensuring that the effective fine-tuned weights will remain in the column space of the pre-trained weight matrix. We evaluate both standard fine-tuning performance (effectiveness) and the robustness of VeFA against LoRA on multiple tasks, including image classification, NLU, and NLG benchmarks. Across these tasks, VeFA achieves comparable fine-tuning performance while consistently showing less degree of forgetting.
@@ 3.1 Problem Setup
We evaluate the robustness/forgetting of fine-tuning methods using established criteria from the literature (Wortsman et al., 2022): fine-tuning on a single downstream dataset should not reduce accuracy on other downstream datasets compared to the zero-shot model. In this work, robustness is therefore interpreted as being inversely related to forgetting—methods that induce less forgetting are regarded as more robust.
@@ 3.2 Vector-based Feature Space Adaptation (VeFA)
Inspired by the EEM view of lurking variables, we propose to perform fine-tuning directly in the feature space via Vector-based Feature Space Adaptation (VeFA). Concretely, we keep $\boldsymbol{W}_{0}^{(\ell)}$ frozen and introduce a diagonal feature-wise scaling matrix $\boldsymbol{\Lambda}^{(\ell)}$, yielding the following model for the $\ell$-th layer:
@@
To compare weight-space and feature-space adaptation schemes, we first give a conceptual illustration. For weight-space adaptation, fine-tuning treats the neurons as fixed and updates the connection weights between them. Both hidden layers and the output layer acquire new trainable parameters. For feature-space adaptation, the original connection weights are frozen, and only lightweight, neuron-wise transformations are learned on top of the pre-trained features. This keeps the representational backbone intact, while still allowing task-specific adaptation through feature scaling.
To better understand their differences, we then give a simple one-dimensional example. Suppose that the pre-trained model is $y=5x$, and the downstream dataset consists of $\{[0.2,0.4],[0.6,1.1],[1.1,2.2],[1.6,2.8]\}$. Using gradient descent, we estimate $\delta_{1}$ in $y=(5+\delta_{1})x$ (weight-space adaptation) and $\delta_{2}$ in $y=5(x+\delta_{2}x)$ (feature-space adaptation). The final results are shown in the Fig. 3. In weight-space fine-tuning, the model adapts by directly updating the parameters, which shifts the regression line (purple dashed) away from the pre-trained function $f(x)=5x$ in second sub-figure. In feature-space fine-tuning, the pre-trained model is kept frozen and adaptation is achieved by learning a lightweight transformation $\Delta(x)$ applied to the input features (orange dashed line in first sub-figure), resulting in the orange dashed regression line in second sub-figure.
@@ 4 Theoretical Analysis
For VeFA, we further prove that the first loss term in Eq. 9 has an upper bound, as specified by Proposition 4.2 (proof can be seen in Appendix D.2). Notably, the bound depends only on the maximum absolute entry of the diagonal vector $\boldsymbol{\lambda}$, so enforcing $\|\boldsymbol{\lambda}\|_{\infty}\leq\gamma$ yields a global bound on the total forgetting under VeFA. In the next section, we also show with experiments that, for most tasks, $\max_{j}|\lambda_{j}|$ of the fine-tuned matrices remains small even without any explicit constraint. For more challenging tasks, one can additionally impose an explicit upper bound on $\max_{j}|\lambda_{j}|$, which directly controls the overall forgetting (seen in Appendix E.1).
@@
In summary, VeFA introduces no intruder dimensions and therefore avoids the associated intruder-induced forgetting. Within the original main column space of $\boldsymbol{W}_{0}$, the transformation loss is tightly bounded and controlled by the small feature-wise scaling factors $\boldsymbol{\lambda}$, so the overall forgetting under VeFA remains bounded. Consequently, VeFA exhibits substantially reduced catastrophic forgetting and correspondingly strong robustness.
@@ 5.1 Experiment setup
Pre-trained models and downstream tasks. To demonstrate the effectiveness of VeFA, we fine-tune three different pre-trained model on their corresponding downstream datasets. Our goals are twofold: (i) to show that VeFA achieves comparable downstream performance to LoRA while using substantially fewer trainable parameters, and (ii) to examine, through multiple cross-task evaluations, whether VeFA attains better performance than LoRA in terms of reduced forgetting and thus improved robustness.
@@
2. Natural Language Understanding (RoBERTa on GLUE(Liu et al., 2019; Wang et al., 2018)). We evaluate RoBERTa on the GLUE benchmark. Since GLUE requires task-specific classification heads, we do not merge all tasks into a single shared head for robustness evaluation. Instead, for each target task, we keep its own fine-tuned classification head, federated-average (McMahan et al., 2017) the VeFA adapters trained on the other GLUE tasks into one shared backbone, and then evaluate this merged backbone together with the target task’s head on that task, without any further fine-tuning. The resulting performance, compared against the zero-shot RoBERTa baseline on the same task, serves as our robustness and forgetting measure for NLU.
3. Natural Language Generation (GPT-2 on E2E (Radford et al., 2019; Novikova et al., 2017)). We fine-tune GPT-2 on the E2E NLG benchmark, which provides a relatively data-rich downstream setting compared to typical few-shot settings. Here the goal is to assess effectiveness: we test whether VeFA’s constrained, channel-wise adaptation can still match LoRA’s performance when more downstream data are available.
@@ 5.2 Fine-tuning CLIP for Image Classification
We follow the setting of previous work (Zhou et al., 2022; Zanella & Ben Ayed, 2024). We evaluate on seven datasets spanning diverse visual domains—satellite imagery (EuroSAT (Helber et al., 2019)), food (Food101 (Bossard et al., 2014)), pet breeds (OxfordPets (Parkhi et al., 2012)), flowers (Flower102 (Nilsback & Zisserman, 2008)), generic objects (Caltech101 (Fei-Fei et al., 2004)), textures (DTD (Cimpoi et al., 2014)), and human actions (UCF101 (Soomro et al., 2012)). Together, these datasets provide a comprehensive benchmark for visual classification.
We compare VeFA, against LoRA (rank $r=2$) in terms of few-shot learning performance and cross-dataset robustness. In few-shot learning, a “shot” refers to the number of labeled training examples provided per class. Few-shot learning closely matches the setting studied in this paper, where downstream data are much smaller in scale than the pre-training datasets and the goal is to mitigate overfitting and forgetting during fine-tuning. We conduct experiments under 1-shot, 4-shot, and 16-shot learning settings. The results are presented in Tab. 1. For the 1-shot and 4-shot settings, we report only effectiveness, whereas for the 16-shot setting, we report both effectiveness and robustness (see the last two rows of Tab. 1). With only 25% of LoRA’s trainable parameters, VeFA achieves higher average performance in the 1-shot, 4-shot, and 16-shot settings, and consistently shows superior robustness across all seven datasets. These results highlight VeFA’s advantage in low-data settings, enabling effective knowledge transfer while better preserving pre-trained representations.
@@ 5.3 Natural Language Understanding
We evaluate on the General Language Understanding Evaluation (GLUE) benchmark (Wang et al., 2018) using RoBERTa-base and RoBERTa-large (Liu et al., 2019). We compare VeFA against LoRA. Our experiment broadly follows the setting in LoRA (Hu et al., 2022): we adapt the query and value projection matrices in each self-attention block and fully train the task-specific classification head. Unlike (Hu et al., 2022), which employs an auxiliary hyperparameter $\alpha$ to rescale gradients in adapted layers, we use separate learning rates for (i) the classification head and (ii) the adapted layers. Learning rates and training epochs are chosen via hyperparameter tuning; full settings appear in Appendix C. We use batch size 64 for RoBERTa-base and 32 for RoBERTa-large respectively.
In line with Hu et al. (2022), we report the number of trainable parameters in the backbone layers, explicitly excluding the task-specific classification head, which is always trained in the standard way. For each setting we run five trials with different random seeds, take the best epoch in each trial, and report the median over these five results. We also reproduce the original LoRA results under the same experimental setup as in Hu et al. (2022) and observe that our SVD-initialized models achieve nearly identical performance, indicating that the SVD approximation does not affect fine-tuning. To evaluate cross-task robustness, we use only the checkpoint with the median validation performance among the 5 runs. For a given GLUE task, we keep its own classification head fixed and federated-average the adapters trained on the other GLUE tasks into a single shared backbone. The robustness metric is then defined as the difference in the task-specific score between this merged-backbone model and the baseline model that uses the original GPT-2 backbone with the same classification head.
The experimental results of effectiveness are presented in Tab. 2. VeFA achieves performance comparable to LoRA across both models, while requiring an order of magnitude fewer parameters. The cross-task robustness results are shown in Table 3, where VeFA achieves stronger cross-task robustness than LoRA.
Table 2: Performance on GLUE with different fine-tuning methods. We report Matthews correlation for CoLA, Pearson correlation for STS-B, and accuracy for all other tasks; higher is better. Results for all methods except VeFA and LoRA are taken from prior work (Hu et al., 2022; Kopiczko et al., 2023).
@@
[…]
Table 3: Cross-task robustness on GLUE via federated merging of medium-checkpoint adapters.
Type SST-2 MRPC CoLA QNLI RTE STS-B AVG
Base LoRA 15.9 0.0 -2.0 3.3 -0.1 15.8 5.48
VeFA 4.7 0.0 -4.0 5.7 1.1 43.7 8.53
Large LoRA -2.0 0.0 0.0 0.1 -1.1 34.0 5.17
VeFA 11.0 6.8 14.1 8.7 11.2 59.0 18.47
@@ 5.4 Natural Language Generation
We evaluate on the E2E benchmark (Novikova et al., 2017), following the experimental protocol used for LoRA (Hu et al., 2022), and fine-tune GPT-2 Medium and Large (Radford et al., 2019). In this setting, the downstream task is more complex, and the dataset is substantially larger than in our visual classification and NLU experiments. Our goal is to test whether VeFA still retains sufficient expressive power under such a challenging regime. For a fair comparison, we run LoRA and VeFA within the same experimental platform. For LoRA, we use the official hyperparameters from (Hu et al., 2022), while a complete list of the hyperparameters used for VeFA is provided in Appendix C. Moreover, We bound VeFA’s scaling vector by clipping each entry of $\boldsymbol{\lambda}$ to [2,2][-2,2], so that $\|\boldsymbol{\lambda}\|_{\infty}\leq 2$ throughout training, as shown in Appendix E.1.
@@ 6 Conclusion
Our empirical results on vision and language tasks show that VeFA can match or even outperform weight-space fine-tuning approaches such as LoRA while using far fewer trainable parameters. VeFA also demonstrates stronger robustness across datasets. Overall, these results highlight VeFA as an efficient alternative for achieving robustness and strong generalization.
Future research could explore alternative projected-gradient schemes that explicitly constrain the update $\boldsymbol{\Delta}$ to the column space of $\boldsymbol{W}_{0}$ while further reducing in-subspace loss. It could also investigate sparsity-inducing constraints on the scaling vector $\boldsymbol{\Lambda}$, connecting VeFA to sparse and mechanistic interpretability.
Impact Statement
This work proposes VeFA, a feature-space fine-tuning method that improves robustness while substantially reducing the number of trainable parameters compared to common weight-space adaptation methods. By restricting adaptation to the column space of $\boldsymbol{W}_{0}$ and avoiding intruder dimensions, VeFA can be particularly useful in low-data settings, where limited flexibility can improve reliability during deployment. We do not anticipate additional societal impacts beyond those commonly associated with improving the efficiency and robustness of model adaptation.
@@ Appendix D Proof of Theoretical Analysis
Appendix E Additional Numerical Experiments
E.1 Maximum VeFA Scaling Coefficient

### v4

@@ Abstract
Catastrophic forgetting is a well-documented challenge in model fine-tuning, particularly under small-data downstream regimes. Existing parameter-efficient fine-tuning methods operate primarily in weight space, which can cause over-specialization of the model to the limited observations of the downstream data. Recent work reveals that one potential mechanism underlying such forgetting is the introduction of intruder dimensions into the representation space during fine-tuning. To mitigate overwriting of pre-trained knowledge and enhance robustness, we propose Vector-based Feature Space Adaptation (VeFA), a new fine-tuning scheme that operates directly in feature space and avoids generating intruder dimensions. VeFA applies element-wise transformation to individual feature channels, ensuring that the effective fine-tuned weights will remain in the column space of the pre-trained weight matrix. We evaluate both standard fine-tuning performance (effectiveness) and the robustness of VeFA against LoRA on multiple tasks, including image classification, NLU, and NLG benchmarks. Across these tasks, VeFA achieves comparable fine-tuning performance while consistently showing less degree of forgetting.
@@ 3.1 Problem Setup
We evaluate the robustness and forgetting of fine-tuning methods using established criteria from the literature (Wortsman et al., 2022): fine-tuning on a single downstream dataset should not reduce accuracy on other downstream datasets compared to the zero-shot model. In this work, robustness is interpreted as inversely related to forgetting: more robust methods are viewed as inducing less forgetting.
@@ 3.2 Vector-based Feature Space Adaptation (VeFA)
Inspired by the EEM view of lurking variables, we propose to perform fine-tuning directly in the feature space via one simple instantiation, Vector-based Feature Space Adaptation (VeFA). Concretely, we keep $\boldsymbol{W}_{0}^{(\ell)}$ frozen and introduce a diagonal feature-wise scaling matrix $\boldsymbol{\Lambda}^{(\ell)}$, yielding the following model for the $\ell$-th layer:
@@
To compare weight-space and feature-space adaptation schemes, we first give a conceptual illustration. For weight-space adaptation, fine-tuning treats the neurons as fixed and updates the connection weights between them. Both hidden layers and the output layer acquire new trainable parameters. For feature-space adaptation, the original connection weights are frozen, and only lightweight, neuron-wise transformations are learned on top of the pre-trained features. This keeps the representational backbone intact, while still allowing task-specific adaptation through feature scaling. We also provide one numerical example on Appendix F.1.
@@ 4 Theoretical Analysis
For VeFA, we further prove that the first loss term in Eq. 9 has an upper bound, as specified by Proposition 4.2 (proof can be seen in Appendix D.2). Notably, the bound depends only on the maximum absolute entry of the diagonal vector $\boldsymbol{\lambda}$, so enforcing $\|\boldsymbol{\lambda}\|_{\infty}\leq\gamma$ yields a global bound on the total forgetting under VeFA. In the next section, we also show with experiments that, for most tasks, $\max_{j}|\lambda_{j}|$ of the fine-tuned matrices remains small even without any explicit constraint. For more challenging tasks, one can additionally impose an explicit upper bound on $\max_{j}|\lambda_{j}|$, which directly controls the overall forgetting (seen in Appendix F.6).
@@
In summary, VeFA introduces no intruder dimensions and, therefore, avoids the associated intruder-induced forgetting. Within the original main column space of $\boldsymbol{W}_{0}$, the transformation loss is tightly bounded and controlled by the small feature-wise scaling factors $\boldsymbol{\lambda}$, so the overall forgetting under VeFA remains bounded. Consequently, VeFA exhibits substantially reduced catastrophic forgetting and, correspondingly, strong robustness.
@@ 5.1 Experiment setup
Pre-trained models and downstream tasks. To demonstrate the effectiveness of VeFA, we fine-tune three different pre-trained models on their corresponding downstream datasets. Our goals are twofold: (i) to show that VeFA achieves downstream performance comparable to LoRA while using substantially fewer trainable parameters, and (ii) to examine, through multiple cross-task evaluations, whether VeFA attains better performance than LoRA in terms of reduced forgetting, and thus improved robustness.
@@
2. Natural Language Understanding (RoBERTa on GLUE (Liu et al., 2019; Wang et al., 2018)). We evaluate RoBERTa on the GLUE benchmark. Since GLUE uses task-specific classification heads, we consider two complementary evaluation settings for cross-task robustness. (1) Pairwise cross-task evaluation. For each source task, we fine-tune an adapter on its training set and evaluate the resulting model on the remaining GLUE tasks. (2) Merge-based cross-task evaluation. For each target task, we keep its fine-tuned head fixed and merge adapters trained on the other tasks into a shared backbone. By default, we use federated averaging (FedAvg) (McMahan et al., 2017). Since GLUE tasks may exhibit non-trivial correlations, we further stress-test the merging procedure by sampling the merging coefficients from a Dirichlet distribution with hyperparameter $\tau$ and forming a convex combination of adapters (Chen & Kwok, 2024). The resulting performance, compared to the zero-shot RoBERTa baseline on the same task, serves as our robustness measure for NLU.
3. Natural Language Generation (GPT-2 on E2E (Radford et al., 2019; Novikova et al., 2017)). We fine-tune GPT-2 on the E2E NLG benchmark, which provides a relatively data-rich downstream setting compared to typical few-shot settings. Here, the goal is to assess effectiveness: we test whether VeFA’s constrained, channel-wise adaptation can still match LoRA’s performance when more downstream data are available.
@@ 5.2 Fine-tuning CLIP for Image Classification
We follow the setting of previous work (Zhou et al., 2022b; Zanella & Ben Ayed, 2024). We evaluated seven datasets spanning diverse visual domains—satellite imagery (EuroSAT (Helber et al., 2019)), food (Food101 (Bossard et al., 2014)), pet breeds (OxfordPets (Parkhi et al., 2012)), flowers (Flower102 (Nilsback & Zisserman, 2008)), generic objects (Caltech101 (Fei-Fei et al., 2004)), textures (DTD (Cimpoi et al., 2014)), and human actions (UCF101 (Soomro et al., 2012)). Together, these datasets provide a comprehensive benchmark for visual classification.
We compare VeFA with prompt-based methods, adapter-based methods, and LoRA (rank $r=2$) in terms of few-shot performance, using a total of eight baselines. We use $r=2$ to match the setting in previous work (Zanella & Ben Ayed, 2024), where it is shown to be sufficient for a few-shot adaptation. In few-shot learning, a shot refers to the number of labeled training examples provided per class. We conducted experiments under 1-shot, 4-shot, and 16-shot learning settings. The results for VeFA and LoRA are reported in Tab. 1. A complete set of results on effectiveness and robustness is provided in Appendices F.2 and F.3, respectively. For the 1-shot and 4-shot settings, we report only effectiveness, whereas for the 16-shot setting, we report both effectiveness and robustness (see the last two rows of Tab. 1). With only 25% of LoRA’s trainable parameters, VeFA achieves higher average performance in the 1-shot, 4-shot, and 16-shot settings and consistently demonstrates stronger robustness across all seven datasets. In addition, VeFA achieves the best accuracy among all competing baselines. These results highlight VeFA’s advantage in low-data settings, enabling effective knowledge transfer while better preserving pre-trained representations.
@@ 5.3 Natural Language Understanding
We evaluated on the General Language Understanding Evaluation (GLUE) benchmark (Wang et al., 2018) using RoBERTa-base and RoBERTa-large (Liu et al., 2019). We compare VeFA against LoRA and six other PEFT methods. Our experiment broadly follows the setting in LoRA (Hu et al., 2022): we adapt each self-attention block and fully train the task-specific classification head. Unlike (Hu et al., 2022), which employs an auxiliary hyperparameter $\alpha$ to rescale gradients in adapted layers, we use separate learning rates for (i) the classification head and (ii) the adapted layers. Learning rates and training epochs are chosen via hyperparameter tuning; full settings appear in Appendix C. We use batch size 64 for RoBERTa-base and 32 for RoBERTa-large.
In line with Hu et al. (2022), we report the number of trainable parameters in the backbone layers, explicitly excluding the task-specific classification head. For each setting, we run five trials with different random seeds, take the best epoch in each trial, and report the median over these five results. We also reproduce the original LoRA results under the same experimental setup as in Hu et al. (2022) and observe that our SVD-initialized models achieve nearly identical performance, indicating that the SVD approximation does not affect fine-tuning. To evaluate cross-task robustness, we use the checkpoint with the median validation performance among the five runs. For model merging, we consider both federated averaging and Dirichlet-based merging with $\tau=10$ (near-uniform) and $\tau=1$ (diverse weighting). For each setting, we draw five independent samples of the merging coefficients.
The key results are summarized in Tab. 2. For completeness, Appendix F.4 reports all effectiveness results, and Appendix F.5 reports all robustness results, including pairwise cross-task evaluation and different merging strategies. In terms of effectiveness, VeFA achieves performance comparable to LoRA and other PEFT methods across both backbones while using substantially fewer trainable parameters. Moreover, across all cross-task robustness settings, VeFA-based adapters show better compatibility and consistently outperform LoRA. These results indicate that VeFA provides stronger cross-task robustness than LoRA.
Table 2: Accuracy and Robustness on the GLUE Benchmark. We report Matthews correlation for CoLA, Pearson correlation for STS-B, and accuracy for all other tasks; higher is better. Results for all methods except VeFA and LoRA are taken from prior work (Hu et al., 2022; Kopiczko et al., 2023; Zhang et al., 2023a).
@@ 5.4 Natural Language Generation
We evaluated on the E2E benchmark (Novikova et al., 2017), following the experimental protocol used for LoRA (Hu et al., 2022), and fine-tuned GPT-2 Medium and Large (Radford et al., 2019). In this setting, the downstream task is more complex and the dataset is substantially larger than in our visual classification and NLU experiments. Our goal is to test whether VeFA still retains sufficient expressive power under such a challenging regime. For a fair comparison, we run LoRA and VeFA within the same experimental platform. For LoRA, we use the official hyperparameters from (Hu et al., 2022), while a complete list of the hyperparameters used for VeFA is provided in Appendix C. Moreover, we bound VeFA’s scaling vector by clipping each entry of $\boldsymbol{\lambda}$ to [2,2][-2,2], so that $\|\boldsymbol{\lambda}\|_{\infty}\leq 2$ throughout training.
@@
5.5 Ablation Study and Further Analysis
We conduct an ablation study to test whether intruder dimensions are truly related to degraded cross-task robustness. Concretely, for the cross-task LoRA models, we apply a projection $P_{\mathcal{S}_{0}}$ that constrains the low-rank update to lie in the pretrained weight subspace $\mathcal{S}_{0}$. The results are shown in Fig. 3. Projection often yields recovery, rescuing a nontrivial portion of harmful-transfer points, which supports our hypothesis that intruder directions contribute to reduced cross-task robustness.
Additional ablation studies on $\boldsymbol{\lambda}$-clipping and the magnitude of adaptation are provided in Appendix E. A further analysis of intruder dimensions in LoRA is provided in Appendix F.7, where we show that the intruder component of $\boldsymbol{\Delta}$ is often non-negligible in magnitude across layers.
Figure 3: Projection mitigates harmful transfer
@@ 6 Conclusion
Our empirical results on vision and language tasks show that VeFA can match or even outperform weight-space fine-tuning approaches while using far fewer trainable parameters. VeFA also demonstrates stronger robustness across datasets. Future research could explore alternative projected-gradient schemes that constrain the update $\boldsymbol{\Delta}$ to $\mathcal{S}_{0}$ while further reducing in-subspace loss. It could also investigate sparsity-inducing constraints on the scaling vector $\boldsymbol{\Lambda}$, connecting VeFA to sparse and mechanistic interpretability.
@@ Appendix D Proof of Theoretical Analysis
Appendix E Additional Ablation Study
Ablation on $\lambda$-Clipping. We first examine the effect of $\lambda$-clipping on VeFA’s performance. On the E2E benchmark, we consider two ablation settings: VeFA with explicit clipping and VeFA without clipping. The results are reported in Tab. 6. Overall, the two settings yield comparable performance, indicating that clipping does not noticeably reduce expressive power, even on more complex tasks.
Table 6: Performance Comparison of VeFA With and Without Clipping on the E2E Benchmark
Type # Params BLEU NIST MET ROUGE-L CIDEr
Med VeFA (w/o clipping) 0.073M 66.42 8.5852 44.40 66.46 2.2134
VeFA (w/ clipping) 0.073M 66.35 8.5628 44.79 66.93 2.2274
Large VeFA (w/o clipping) 0.14M 67.62 8.5956 46.04 68.77 2.3858
VeFA (w/ clipping) 0.14M 67.38 8.5889 46.11 68.86 2.3684
Magnitude of Adaptation. In Fig. 5, we visualize the layer-wise magnitude of the learned $\boldsymbol{\lambda}$ vectors after fine-tuning on RTE and MRPC. Overall, the largest adaptation happens for query matrices compared to the value ones, indicating a larger need or ease for finetuning a model there. Moreover, consistent with observations in prior parameter-efficient adaptation studies (Kopiczko et al., 2023; Zhang et al., 2023b; Liu et al., 2022), we observe stronger adaptations in later layers than in earlier layers.
(a) RTE
(b) MRPC
Figure 5: Magnitude of the adapted $\boldsymbol{\lambda}$ vector for query and value matrices across layers for RoBERTa-L.
Appendix F Additional Numerical Experiments
F.1 Comparison of Weight-Space and Feature-Space Adaptation
To better understand the differences between feature, we then give a simple one-dimensional example. Suppose that the pre-trained model is $y=5x$, and the downstream dataset consists of $\{[0.2,0.4],[0.6,1.1],[1.1,2.2],[1.6,2.8]\}$. Using gradient descent, we estimate $\delta_{1}$ in $y=(5+\delta_{1})x$ (weight-space adaptation) and $\delta_{2}$ in $y=5(x+\delta_{2}x)$ (feature-space adaptation). The final results are shown in the Fig. 6. In weight-space fine-tuning, the model adapts by directly updating the parameters, which shifts the regression line (purple dashed) away from the pre-trained function $f(x)=5x$ in second sub-figure. In feature-space fine-tuning, the pre-trained model is kept frozen and adaptation is achieved by learning a lightweight transformation $\Delta(x)$ applied to the input features (orange dashed line in first sub-figure), resulting in the orange dashed regression line in second sub-figure.
F.2 Effectiveness on the Visual Classification Benchmark
Few-shot learning closely matches the setting studied in this paper, where downstream data are much smaller in scale than the pre-training datasets and the goal is to mitigate overfitting and forgetting during fine-tuning. We compare VeFA with several prompt-based methods, including CoOp (4) with four learnable context tokens (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), PLOT++ (Chen et al., 2022), and ProGrad (Zhu et al., 2023) with 16 learnable tokens, which is conceptually related to our approach in that it constrains updates in the few-shot regime. We also report results for adapter-based methods, including TIP-Adapter-F (Zhang et al., 2022), where we reduce the validation set to a reasonable size, TaskRes (Yu et al., 2023) and CLIP-Adapter.
The final comparison is reported in Tab.7. Our VeFA outperforms all eight baselines across seven datasets, validating the effectiveness of VeFA.
Table 7: Performance comparison between VeFA and eight baselines across different datasets under zero-shot, 1-shot, 4-shot, 16-shot fine-tuning. The visual backbone is ViT-B/16. Best results are in bold.
Shots Type Caltech101 Food101 Oxford Pets Oxford Flowers EuroSAT DTD UCF101 Avg
0 CLIP 93.2 85.2 88.2 67.3 42.3 44.3 65.0 69.4
1 CoOp (4) 93.2 82.6 90.3 72.7 50.9 50.1 70.7 72.9
CoCoOp 94.1 84.9 91.9 73.4 55.4 52.6 70.4 74.7
TIP-Adapter-F 94.0 85.8 90.6 83.8 67.8 51.6 73.4 78.1
CLIP-Adapter 92.0 86.1 89.0 71.3 49.3 44.2 66.9 71.3
PLOT++ 94.3 86.2 91.9 80.5 65.4 54.6 74.3 78.2
TaskRes 93.6 84.6 90.2 81.7 65.4 53.8 71.7 77.3
ProGrad 93.5 84.9 91.4 80.9 57.0 52.8 73.3 76.3
LoRA 94.0 82.3 91.8 79.4 71.8 53.0 75.4 78.2
VeFA (Ours) 94.1 86.3 93.3 84.3 73.3 55.0 74.2 80.1
4 CoOp (4) 94.5 83.5 92.3 86.6 65.8 58.5 78.1 79.9
CoCoOp 94.8 86.3 92.7 81.5 61.7 55.7 75.3 78.3
TIP-Adapter-F 94.8 86.5 91.9 92.1 76.8 59.8 78.1 82.9
CLIP-Adapter 94.0 86.5 90.8 73.1 51.2 46.1 70.6 73.2
PLOT++ 95.1 86.5 92.6 92.9 83.2 62.4 79.8 84.6
TaskRes 95.0 86.0 91.9 85.0 74.2 60.1 76.2 81.2
ProGrad 94.4 85.4 92.1 91.1 69.6 59.7 77.9 81.5
LoRA 95.2 81.6 90.8 93.8 87.5 62.2 78.5 84.2
VeFA (Ours) 95.6 86.0 93.4 93.0 87.1 65.7 80.2 85.7
16 CoOp (4) 95.5 85.1 92.4 96.4 83.5 69.2 81.9 86.3
CoCoOp 95.1 87.4 93.4 89.1 73.6 63.7 77.2 82.3
TIP-Adapter-F 95.7 86.8 92.6 96.2 85.9 70.8 83.9 87.4
CLIP-Adapter 94.9 87.1 92.3 92.9 71.4 59.4 80.2 82.6
PLOT++ 96.0 87.1 93.6 97.6 92.0 71.4 85.3 89.0
TaskRes 95.8 86.9 92.4 97.5 82.7 71.5 84.0 87.3
ProGrad 95.9 85.8 92.8 96.6 83.6 68.8 82.7 86.6
LoRA 96.3 84.0 91.3 98.2 90.2 71.8 86.2 88.3
VeFA (Ours) 96.4 87.8 94.4 97.5 91.3 72.5 86.4 89.5
F.3 Robustness on the Visual Classification Benchmark
Figure 7 shows a clear difference in cross-dataset transfer behavior between VeFA and LoRA. VeFA exhibits much more stable transfer, with most entries staying close to zero and several directions even showing small positive gains, indicating that adaptation causes limited disruption to the pre-trained representations. In contrast, LoRA is dominated by negative shifts, with many strongly negative cells and several severe drops (e.g., when adapting on EuroSAT), suggesting a higher risk of harmful transfer and forgetting. Overall, VeFA consistently reduces both the frequency and magnitude of performance degradation across datasets, demonstrating more reliable and robust cross-domain generalization.
(a) VeFA
(b) LoRA
Figure 7: Cross-dataset transfer heatmaps. Each entry shows the performance change relative to zero-shot learning (diagonal entries are set to 0).
F.4 Effectiveness on the GLUE Benchmark
A broad comparison of VeFA against full fine-tuning, BitFit, adapter-based methods, LoRA, and its variants is shown in Tab.8. VeFA achieves comparable performance—and often outperforms many alternatives—while using the fewest trainable parameters.
Table 8: Performance on GLUE with different fine-tuning methods. We report Matthews correlation for CoLA, Pearson correlation for STS-B, and accuracy for all other tasks; higher is better. Results for all methods except VeFA and LoRA are taken from prior work (Hu et al., 2022; Kopiczko et al., 2023; Zhang et al., 2023a).
Type # Params SST-2 MRPC CoLA QNLI RTE STS-B AVG
Base FT 125M 94.8 90.2 63.6 92.8 78.7 91.2 85.2
BitFit 0.1M 93.7 92.7 62.0 91.8 81.5 90.8 85.4
[…]
LoRA-FA 1.8M 94.8 90.0 63.6 92.5 67.9 89.6 83.1
[…]
Large FT 355M 96.4 90.9 68.0 94.7 86.6 92.4 88.1
[…]
LoRA-FA 3.7M 96.0 90.0 68.0 94.4 86.1 92.0 87.7
[…]
F.5 Robustness on the GLUE Benchmark
Table 9 summarizes cross-task robustness on GLUE under both RoBERTa-base (B) and RoBERTa-large (L). Overall, VeFA consistently achieves markedly higher robustness than LoRA across all robustness settings, while using around $16\times$ fewer trainable parameters (0.018M vs. 0.3M). On RoBERTa-base, VeFA improves the average robustness from 4.0 to 18.8 for $\tau=1$ and from 4.3 to 17.2 for $\tau=10$. It also outperforms LoRA under $\bar{R}$ (6.5 to 11.3) and FedAvg (5.48 to 8.53). The gains are particularly pronounced on STS-B (e.g., 44.9–46.9 for VeFA vs. 6.1–9.5 for LoRA), indicating substantially better transferability across tasks. The trend is even clearer on RoBERTa-large, where LoRA’s average robustness is close to zero or even negative in several settings (e.g., 1.5 and 1.3 for $\tau\in\{1,10\}$), while VeFA maintains strong positive robustness (about 20.0 and 19.6). VeFA also remains consistently strong under the other merging strategies on RoBERTa-large (e.g., 18.47 with FedAvg and 18.7 with $\bar{R}$), suggesting that feature-space adaptation yields adapters that are more compatible for cross-task reuse and combination than weight-space LoRA updates.
Table 9: Robustness on GLUE benchmark. We report Matthews correlation for CoLA, Pearson correlation for STS-B, and accuracy for all other tasks; higher is better.
Type # Params SST-2 MRPC CoLA QNLI RTE STS-B AVG
[…]
LoRA (FedAvg) 0.3 M 15.9 0.0 -2.0 3.3 -0.1 15.8 5.48
VeFA (FedAvg) 0.018 M 4.7 0.0 -4.0 5.7 1.1 43.7 8.53
LoRA ($\bar{R}$) 0.3 M 7.16 18.2 4.9 -4.1 4.3 8.6 6.5
VeFA ($\bar{R}$) 0.018 M 13.38 9.3 -1.6 4.4 11.2 31.2 11.3
[…]
LoRA (FedAvg) 0.3 M -2.0 0.0 0.0 0.1 -1.1 34.0 5.17
VeFA (FedAvg) 0.018 M 11.0 6.8 14.1 8.7 11.2 59.0 18.47
LoRA ($\bar{R}$) 0.3 M 2.26 12.1 4.8 3.8 4.1 18.2 7.5
VeFA ($\bar{R}$) 0.018 M 8.90 23.4 2.6 9.1 12.8 55.3 18.7
F.6 Maximum VeFA Scaling Coefficient